OpenStudy (anonymous):

Solve the system of equations. Enter your answer as an ordered pair. 65x - 7y = 371 -11x + 7y = 7

OpenStudy (chaise):

65x - 7y = 371 -11x + 7y = 7 Plus the two equations together to eliminate the y. Solve for x Substitute your x value back into one of your equations. Solve for y. Write as: (x,y)